# Causal structure and topology change in (2+1)-dimensional simplicial gravity

### v1

@@ 3.1 Local causality conditions in Lorentzian triangulation
In a generic simplicial Lorentzian geometry, causality can be meaningfully assigned to sub-simplices of co-dimension two or higher [35, 12]. This ensures that the surrounding geometry admits a nontrivial causal structure when projected orthogonally to the sub-simplex. For instance in 1+1 dimensions, where triangles are the fundamental building blocks, only vertices (co-dimension 2) admit a causal structure. In 2+1 dimensions, both edges (co-dimension 2) and vertices (co-dimension 3) support well-defined causality conditions.
Consider a vertex $v$ in the bulk of a $d$-dimensional triangulation. Its causal structure is determined by examining the collection of all $d$-simplices that share the vertex. The star of $v$, denoted by ${\rm St}(v)$ is defined as the union of all the simplices that are incident on $v$. Topologically, ${\rm St}(v)$ is homeomorphic to a closed $d$-ball. Its boundary $\partial{\rm St}(v)$ is given by the union of all $(d-1)$–simplices that are opposite to $v$. The boundary star also forms a triangulation of a sphere $S^{d-1}$. The causality condition at the vertex $v$ is then inferred by intersecting the lightcones emanating from the vertex with the boundary $\partial{\rm St}(v)$. We will make this construction explicit in the (2+1)–dimensional case. The vertex is said to be causally regular if this intersection yields exactly two disconnected components, corresponding to past and future lightcones. Configurations that deviate from this canonical structure are referred to as causally irregular. These irregular causal structures typically appear in pathological spacetimes such as those that describe spatial topology change or closed timelike curves.
@@
More broadly, in any dimension $d$, one can define causality conditions not only at vertices but also at edges, triangles, and other faces (up to hinges of co-dimension two). These conditions, referred to as vertex–causality, edge–causality, triangle–causality, and so on, generalizes the notion of vertex–causality to higher-dimensional subsimplices. For any sub-simplex $\sigma_{k}$ of dimension $k\leq d-2$, its causal structure is determined by projecting $\sigma_{k}$ to a point $p$. This projection [^4] The projection of a $k$-simplex $\sigma_{k}$ leads to a quotient space $\mathbb{R}^{1,d-1}/V_{k}$, where $k<d$ and $V_{k}$ is a subspace of $\mathbb{R}^{1,d-1}$ spanned by $\sigma_{k}$. This quotient space is isomorphic to either Euclidean or Minkowski spacetime if $\sigma_{k}$ is timelike or spacelike respectively [11]. yields a simplical complex in the local neighbourhood of the point, ${\rm St}(p)$, which is homeomorphic to a ball in $\mathbb{R}^{d-k}$. The causal structure of the sub-simplex $\sigma_{k}$ is then determined by the intersection of the lightcones emanating from the projected point with the boundary of this spherical ball. In general, projecting out a hinge always reduces to a two dimensional local neighbourhood. Hence, ‘hinge–causality’ becomes equivalent to causal structure at vertices in two dimensional spacetimes, which is a well-understood concept in simplicial gravity [12, 11].
@@
In this article, we focus on the analysis of vertex–causality in 2+1 dimensions. Our goal is to construct a practical framework for determining vertex causal structure using local data (specifically, squared edge lengths) and to assess its implications for discrete curvature, topology change, and quantum gravitational path integrals. To this end, we classify the causal types of triangles and tetrahedra based on edge signatures and develop a practical procedure for reconstructing the lightcone configurations at a vertex from purely local data (edge lengths of neighbouring simplices).
@@ 3.2 Classification of triangles in Minkowski spacetime
$\Theta_{v_{0}}=-\imath\log_{-}\left(\frac{(s_{1}+s_{2}-s_{12})-\imath\sqrt{4s_{1}s_{2}-(s_{1}+s_{2}-s_{12})^{2}}}{2\sqrt{s_{1}}\sqrt{s_{2}}}\right)\,,\quad$ (8)
@@ 4.1 Causal classifications of Lorentzian tetrahedra
Given the squared edge lengths of a tetrahedron, its embedding in $\mathbb{R}^{1,2}$ can be constructed by fixing the reference vertex $v_{0}$ at the origin and determining the exact positions of the other vertices through the edge vectors. These edge vectors, forming the columns of a $3\times 3$ matrix $L$, can be obtained via a decomposition of the Gram matrix of inner products in Eq. (2) as $G(s)=L\eta L^{\top}$, where $\eta={\rm diag}(-1,1,1)$. If the Gram matrix $G(s)$ is non-degenerate and the squared edge lengths satisfy the generalized Lorentzian simplex inequalities (see Proposition 1), then the matrix of edge vectors $L$ can be determined up to global Lorentz transformations. This construction gives an explicit embedding of a tetrahedron in $\mathbb{R}^{1,2}$ and allows us to characterize the causal structure of the tetrahedron relative to the lightcones at $v_{0}$.
To systematically classify the causal types of a tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$, we examine the intersection between the lightcone at $v_{0}$ and the opposite triangle $\Delta(v_{1}v_{2}v_{3})$. Using barycentric coordinates, any point on this triangle can be expressed as
@@
where each point lies in the affine span of its vertices $v_{1},v_{2},v_{3}$, and its causal character is determined by the Minkowski norm $\langle x(\alpha,\beta),x(\alpha,\beta)\rangle$. The set of points where this norm vanishes (where the opposite triangle intersects the lightcone) traces out conic sections [^5] The conic sections may appear as hyperbola, parabola or an ellipse on the triangle depending on the causal structure. which divides the triangle into spacelike and timelike regions. The vertex coordinates used in these constructions are not chosen externally but are implicitly determined by the intrinsic edge length data via the Gram matrix. Consequently, all geometric and causal information, including lightcone intersections and causal types, are derived from the edge lengths alone, without reference to any background coordinates.
@@ A. Timelike vertices:
For type A1, the opposite triangle does not intersect the lightcone and lies within one timelike region. In the type A2, the triangle intersects both the past and future lightcones of $v_{0}$, resulting in two null curves which separate the triangle into two disconnected timelike regions and one spacelike region.
@@ B. Two timelike, one spacelike vertices:
These sets of length data satisfy the Lorentzian simplex inequalities. In the type B1, where both timelike vertices lie in the same causal region, the intersection of the lightcone with the opposite triangle produces a single null curve, dividing the triangle into one connected timelike and one connected spacelike region. In contrast, the type B2 tetrahedra has the two timelike vertices in opposite causal regions. The corresponding lightcone intersection consists of two null curves, separating the triangle into two disconnected timelike regions and one spacelike region as shown in Figure 4.
@@ C. One timelike, two spacelike vertices:
  • C3. Two-lightcone instersections: Here, unlike in the type C2 above, the second lightcone intersects the boundary edge $e(v_{2}v_{3})$. As a result, the triangle $\Delta(v_{1}v_{2}v_{3})$ intersects both future and past lightcones incident at $v_{0}$. The triangle $\Delta(v_{0}v_{2}v_{3})$ is again of type T5.
@@
${\cal S}_{\bf C1}=(-1,1,1,-1,1,1),\quad{\cal S}_{\bf C2}=(-2,2,2,2,2,9),\quad{\cal S}_{\bf C3}=(-1,1,1,-1,-6,6)\,,\quad$
@@ D. Three spacelike vertices:
  • D2. One-lightcone intersection inside triangle: The interior of the triangle $\Delta(v_{1}v_{2}v_{3})$ intersects a exactly one lightcone, but the boundary edges do not. The resulting intersection gives a closed null curve inside the opposite triangle. All edges of the triangle are spacelike [^7] For the tetrahedron type D2, we only encountered examples or configurations where all the triangular faces of the tetrahedron are spacelike, i.e., of type E..
@@
Consider a tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$ in Minkowski spacetime $\mathbb{R}^{1,2}$ with no null edges and null triangular faces. The causal structure at a reference vertex $v_{0}$ (placed at the origin) is uniquely determined by the causal positions of the vertices $(v_{1},v_{2},v_{3})$ relative to the two lightcones at $v_{0}$ and the intersection of the lightcones with the opposite triangle $\Delta(v_{1}v_{2}v_{3})$. There are exactly 13 distinct vertex causal types of Lorentzian tetrahedra under these conditions.
@@ 4.2 Causal structure at a vertex in a triangulation
A key requirement in constructing a well-defined causal structure in a Lorentzian triangulation is the continuity of null curves (arising from lightcone intersections) across shared faces of adjacent simplices. In Regge calculus, the geometry (provided by the edge lengths) along the interface between a pair of simplices is locally consistent, i.e., the induced metric on a face shared by the two simplices agrees. Thus, one can introduce a coordinate system for any two simplices in which the metric is differentiable across their shared face. Let $\tau_{1}$ and $\tau_{2}$ be two adjacent tetrahedra in ${\rm St}(v)$ sharing a triangle, and let $C_{1}$ and $C_{2}$ be the planar conic sections induced by the intersection of the lightcone at $v$ with the respective boundary triangles in $\partial{\rm St}(v)$. The adjacent opposite triangles share an edge $e\in\partial{\rm St}(v)$. Then the condition
@@
ensures that the null curves induced by the lightcones at $v$ intersect the same points along the edge $e$. However, since $C_{1}$ and $C_{2}$ lie in distinct planes, their tangent directions $\frac{dC_{i}}{d\gamma}$ (where $\gamma$ denotes an affine parameter along the null curves) generally differ across the shared edge $e$. As a result, the null curves are not necessarily smooth across the edge. This reflects the piecewise nature of the triangulation and the local curvature concentrated at edges. Together, these conditions ensure that the resulting conic sections are continuous and piecewise smooth [^9] This is one reason why null faces were excluded from the construction so far. This implicitly excludes a class of causal irregularities where two lightcones may intersect tangentially along a shared null ray. on $\partial{\rm St}(v)$. This gluing condition is reminiscent of treatments of geodesic propagation in Regge calculus [65], describing a discrete refraction law for lightrays across hinges. There, geodesics refract across faces according to the local geometry, but do not exhibit singularities or divergences at hinges. Analogously, the smooth propagation of null curves across simplices reflects the local flatness and consistent gluing of the simplicial metric data.
@@
    Use the squared edge length data $\cal S_{X}$ of each tetrahedron to determine its causal type, and trace out the intersection of the lightcones at $v$ with the corresponding opposite triangle in $\partial\mathrm{St}(v)$, which generically yields a conic section.
@@
    Connect the conic sections of neighboring tetrahedra. Identify and count the number of connected spacelike regions $N_{\rm sl}$, and timelike regions $N_{\rm tl}$ on $\partial\mathrm{St}(v)$, separated by the piecewise null curves.
@@
Figure 7: Illustration of a regular causal structure at a bulk vertex. The lightcone intersection patterns results in two timelike regions and one spacelike region separated by null curves on the boundary sphere $S^{2}$.
The resulting data $(N_{\rm sl},N_{\rm tl})$, which enumerate the number of connected spacelike and timelike regions on the boundary spherical triangulation $\partial\mathrm{St}(v)$, provides a precise and geometrically meaningful characterization of the local causal structure around each vertex $v$ in a Lorentzian triangulation. In particular, a vertex $v$ is said to be causally regular iff $(N_{\rm sl},N_{\rm tl})=(1,2)$, i.e., there are exactly two closed null curves which intersect the triangulated boundary sphere as displayed in Figure 7. These two null curves partition $\partial\mathrm{St}(v)$ into a single connected spacelike region and two connected timelike regions, reflecting the canonical causal ordering expected at a generic spacetime point in continuous Lorentzian geometry. This regular configuration mirrors the local causal structure of smooth Lorentzian manifolds, where at each point the lightcones divide the tangent space into precisely two causally disconnected timelike sectors and one spacelike sector in between. Any deviation from this regular structure, where $(N_{\rm sl},N_{\rm tl})\neq(1,2)$, signals a causal irregular structure. Such configurations represent a breakdown of the standard Lorentzian causal structure, and may signal causal pathologies, local signature changes or topology change in discrete geometries.

### v2

@@ 3.3 Local vertex–causality conditions in a triangulation
In a generic simplicial Lorentzian geometry, causality conditions can be meaningfully assigned to subsimplices of codimension–2 or higher [35, 12]. This ensures that the surrounding geometry admits a nontrivial lightcone structure when projected orthogonally to the subsimplex. For instance in 1+1 dimensions, where triangles are the fundamental building blocks, only vertices (codimension–2) admit a causal structure. In 2+1 dimensions, both edges (codimension–2) and vertices (codimension–3) support well-defined causality characterizations.
Consider now a vertex $v$ in the bulk of a $d$-dimensional Lorentzian triangulation. We use the term causal structure of the vertex or vertex–causality to denote the causal structure of its local neighbourhood ${\rm St}(v)$ relative to $v$. The lightcones $\mathcal{N}^{\pm}_{i}(v)$ centered at $v$ within each incident simplex $\sigma_{i}\in{\rm St}(v)$ intersect the corresponding boundary faces, collectively partitioning the boundary triangulation $\partial{\rm St}(v)$ into regions that are spacelike or timelike separated from $v$. The causal structure of ${\rm St}(v)$ can therefore be inferred from the pattern of the lightcone intersections incident on $v$. We shall make this construction explicit in the (2+1)–dimensional case. A vertex $v$ is said to be causally regular if the combined intersections $\bigcup_{i}\big(\mathcal{N}_{i}^{\pm}(v)\cap\partial\sigma^{d}_{i}\big)$ form exactly two disconnected timelike components separated by one connected spacelike component on $\partial{\rm St}(v)$ relative to $v$. The two timelike disconnected components correspond to the past and future lightcones at $v$. Configurations that deviate from this canonical structure are termed causally irregular. Such irregular causal structures typically arise in pathological spacetimes, for instance those describing spatial topology change.
@@
More generally, in any dimension $d$, one can define causality conditions not only at vertices but also at edges, triangles, and other faces (up to hinges of codimension 2). These conditions, referred to as vertex–causality, edge–causality, triangle–causality, and so on, each describe the causal structure of the local neighbourhood of a subsimplex. Let $\sigma^{k}$ be a Euclidean subsimplex of dimension $k\leq d-2$. Projecting [^4] The projection of a $k$-simplex $\sigma_{k}$ leads to a quotient space $\mathbb{R}^{1,d-1}/V_{k}$, where $k<d$ and $V_{k}$ is a subspace of $\mathbb{R}^{1,d-1}$ spanned by $\sigma_{k}$. This quotient space is isomorphic to either Euclidean or Minkowski spacetime if $\sigma_{k}$ is timelike or spacelike respectively [11]. $\sigma^{k}$ orthogonally to a point $p$, and applying this projection to all simplices in its star ${\rm St}(\sigma^{k})$, yields a lower-dimensional simplicial complex ${\rm St}(p)$ in the local neighbourhood of the point. The resulting complex is homeomorphic to a ball in $\mathbb{R}^{d-k}$. The causal structure of the subsimplex $\sigma^{k}$ is then determined by the intersection of the lightcones emanating from the projected point with the boundary of this spherical ball. In particular, projecting out a hinge always reduces to a two dimensional local neighbourhood. Hence, hinge–causality becomes equivalent to the notion of vertex–causality in two dimensions, which is a well-understood concept in simplicial gravity [12, 11].
@@
In this article, we focus on the analysis of vertex–causality in 2+1 dimensions. Our goal is to construct a practical framework for determining vertex causal structure using local data (specifically, squared edge lengths) and to assess its implications for discrete curvature, topology change, and quantum gravitational path integrals. To this end, we begin by classifying the causal types of triangles and tetrahedra based on edge signatures and develop a practical procedure for reconstructing the lightcone structures at a vertex from purely local data (edge lengths of neighbouring simplices).
@@ 3.4 Classification of triangles in Minkowski spacetime
$\Theta_{v_{0}}=-\imath\log_{-}\left(\frac{(s_{1}+s_{2}-s_{12})-\imath\sqrt{4s_{1}s_{2}-(s_{1}+s_{2}-s_{12})^{2}}}{2\sqrt{s_{1}}\sqrt{s_{2}}}\right)\,$$\quad$ (11)
@@ 4.1 Causal classifications of Lorentzian tetrahedra
Given the squared edge lengths of a tetrahedron, its embedding in $\mathbb{R}^{1,2}$ can be constructed by fixing the reference vertex $v_{0}$ at the origin and determining the exact positions of the other vertices through the edge vectors. These edge vectors, forming the columns of a $3\times 3$ matrix $L$, can be obtained via a decomposition of the Gram matrix of inner products in Eq. (2) as $G(s)=L\eta L^{\top}$, where $\eta={\rm diag}(-1,1,1)$. If the Gram matrix $G(s)$ is non-degenerate and the squared edge lengths satisfy the generalized Lorentzian simplex inequalities (see Proposition 1), then the matrix of edge vectors $L$ can be determined up to global Lorentz transformations. This construction gives an explicit embedding of a tetrahedron in $\mathbb{R}^{1,2}$ and allows us to characterize its causal structure of relative to a vertex (see Definition 2).
To systematically classify the causal types of a tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$, we examine the intersection of the lightcones $\mathcal{N}^{\pm}(v_{0})$ centered at $v_{0}$ with the opposite triangle $\Delta(v_{1}v_{2}v_{3})$. Here, the term ‘causal type’ distinguishes distinct configurations of how the vertices of $\Delta(v_{1}v_{2}v_{3})$ lie relative to the lightcones at $v_{0}$. This information will be essential in the analysis of larger triangulations. Using barycentric coordinates, any point on this triangle can be expressed as
@@
where each point lies in the affine span of its vertices $v_{1},v_{2},v_{3}$. The causal character of a point $x(\alpha,\beta)$ relative to $v_{0}$ is determined by the Minkowski norm $\langle x(\alpha,\beta),x(\alpha,\beta)\rangle$. The set of points where this norm vanishes (where the opposite triangle intersects the lightcone) traces out conic sections [^5] The conic sections may appear as hyperbola, parabola or an ellipse on the triangle depending on the edge lengths of the tetrahedron. which divides the triangle into regions that are spacelike and timelike with respect to $v_{0}$. The coordinates used in these constructions are not chosen externally but are implicitly determined by the intrinsic edge length data via the Gram matrix. Consequently, all geometric and causal information, including lightcone intersections and causal types, are derived from the edge lengths alone, without reference to any background coordinates.
@@ A. Timelike vertices:
For type A1, the opposite triangle does not intersect the lightcone and lies within one timelike region. In the type A2, the triangle intersects both the past and future lightcones of $v_{0}$, resulting in two lightcone intersection curves which separate the triangle into two disconnected timelike regions and one spacelike region.
@@ B. Two timelike, one spacelike vertices:
These sets of length data satisfy the Lorentzian simplex inequalities. In the type B1, where both timelike vertices lie in the same causal region, the intersection of the lightcone with the opposite triangle produces a conic section, dividing the triangle into one connected timelike and one connected spacelike region. In contrast, the type B2 tetrahedra has the two timelike vertices in opposite causal regions. The corresponding lightcone intersection consists of two lightcone intersection curves, separating the triangle into two disconnected timelike regions and one spacelike region as shown in Figure 4.
@@ C. One timelike, two spacelike vertices:
  • C3. Two-lightcone intersections: Here, unlike in the type C2 above, the second lightcone intersects the boundary edge $e(v_{2}v_{3})$. As a result, the triangle $\Delta(v_{1}v_{2}v_{3})$ intersects both future and past lightcones incident at $v_{0}$. The triangle $\Delta(v_{0}v_{2}v_{3})$ is again of type T5.
@@
${\cal S}_{\bf C1}=(-1,1,1,-1,1,1),\quad{\cal S}_{\bf C2}=(-2,2,2,2,2,9),\quad{\cal S}_{\bf C3}=(-1,1,1,-1,-6,6)\,$$\quad$
@@ D. Three spacelike vertices:
  • D2. One-lightcone intersection inside triangle: The interior of the triangle $\Delta(v_{1}v_{2}v_{3})$ intersects a exactly one lightcone, but the boundary edges do not. The resulting intersection gives a closed conic section inside the opposite triangle. All edges of the triangle are spacelike [^7] For the tetrahedron type D2, we only encountered examples or configurations where all the triangular faces of the tetrahedron are spacelike, i.e., of type E..
@@
Consider a tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$ in Minkowski spacetime $\mathbb{R}^{1,2}$ with no null edges and no null triangular faces. The causal structure of $\Delta(v_{0}v_{1}v_{2}v_{3})$ relative to a reference vertex $v_{0}$ (as provided in Definition 2) is uniquely determined by the causal positions of the vertices $(v_{1},v_{2},v_{3})$ relative to the two lightcones at $v_{0}$ and the intersection of the lightcones at $v_{0}$ with the opposite triangle $\Delta(v_{1}v_{2}v_{3})$. There are exactly 13 distinct vertex causal types of Lorentzian tetrahedra under these conditions.
@@ 4.2 Causal structure at a vertex in a triangulation
A key requirement in constructing a well-defined causal structure in a Lorentzian triangulation is the continuity of lightcone intersection curves across shared faces of adjacent simplices. In Regge calculus, the geometry (provided by the edge lengths) along the interface between a pair of simplices is locally consistent, i.e., the induced metric on a face shared by the two simplices agrees. Thus, one can introduce a coordinate system for any two simplices in which the metric is differentiable across their shared face. Let $\tau_{1}$ and $\tau_{2}$ be two adjacent tetrahedra in ${\rm St}(v)$ sharing a triangle, and let $C_{1}$ and $C_{2}$ be the planar conic sections obtained by the intersection of the lightcones at $v$ with the respective boundary triangles in $\partial{\rm St}(v)$. The adjacent opposite triangles share an edge $e\in\partial{\rm St}(v)$. Then the condition
@@
ensures that the lightcone intersection curves are continuous across the shared edge $e$, i.e., they meet at the same set of points along $e$. However, since $C_{1}$ and $C_{2}$ lie in distinct planes, their tangent directions $\frac{dC_{i}}{d\gamma}$ (where $\gamma$ denotes an affine parameter along the null curves) generally differ across the shared edge $e$. As a result, the lightcone intersection curves are not necessarily smooth across the interface. This non-smoothness reflects the piecewise-flat nature of the triangulation and the localized curvature along the edges characteristic of Regge geometry.
Together, these conditions ensure that the resulting conic sections are continuous and piecewise-smooth [^9] This is another reason why null faces were excluded from the construction so far. This implicitly excludes a class of causal irregularities where two lightcones may intersect tangentially along a shared null ray. on $\partial{\rm St}(v)$. This gluing condition is reminiscent of treatments of geodesic propagation in Regge calculus [65], describing a discrete refraction law for lightrays across hinges. There, geodesics refract across faces according to the local geometry without singularities or divergences at the hinges. This reflects the local flatness of each simplex and the consistent gluing of the simplicial metric data across shared faces.
@@
    Use the squared edge length data $\cal S_{X}$ of each tetrahedron to determine its causal type, and trace out the lightcone intersection curves on the corresponding opposite triangles in $\partial\mathrm{St}(v)$, which generically yields conic section(s).
@@
    Connect the conic sections of neighbouring tetrahedra. Identify and count the number of connected spacelike regions $N_{\rm sl}$, and timelike regions $N_{\rm tl}$ on $\partial\mathrm{St}(v)$, separated by the piecewise smooth lightcone intersection curves.
@@
Figure 7: Illustration of a regular causal structure at a bulk vertex. The lightcone intersection patterns results in two timelike regions and one spacelike region separated by lightcone intersection curves on the boundary sphere $S^{2}$.
The resulting data $(N_{\rm sl},N_{\rm tl})$, which enumerate the number of connected spacelike and timelike regions on the boundary spherical triangulation $\partial\mathrm{St}(v)$, provides a precise and geometrically meaningful characterization of the local causal structure around each vertex $v$ in a Lorentzian triangulation. In particular, a vertex $v$ is said to be causally regular iff $(N_{\rm sl},N_{\rm tl})=(1,2)$, i.e., there are exactly two closed lightcone intersection curves which intersect the triangulated boundary sphere as displayed in Figure 7. These two curves partition $\partial\mathrm{St}(v)$ into a single connected spacelike region and two connected timelike regions, reflecting the canonical causal ordering expected at a generic spacetime point in continuous Lorentzian geometry. This regular configuration mirrors the local causal structure of smooth Lorentzian manifolds, where at each point the lightcones divide the tangent space into precisely two causally disconnected timelike sectors and one spacelike sector in between. Any deviation from this regular structure, where $(N_{\rm sl},N_{\rm tl})\neq(1,2)$, signals a causal irregular structure. Such configurations represent a breakdown of the standard Lorentzian causal structure, and may signal causal pathologies, local signature changes or topology change in discrete geometries.